OpenStudy (anonymous):

When 36 is subtracted from the square of a number, the result is five times the number. What is the positive solution?

OpenStudy (anonymous):

x^2 - 36=5x solve it

OpenStudy (anonymous):

can you solve quadratic equation? x^2-5x-36=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

need help ? or already got answer?

OpenStudy (anonymous):

solving it now

OpenStudy (anonymous):

ill wait

OpenStudy (anonymous):

is it 6?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

=(x-9)(x+4)=0 so your positive number is...

OpenStudy (anonymous):

9

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome!